if 10^-5 % react molecules is crossing over the barrier in transition state at 298 K then find out the activation barrier needed to cross the transition state? 

Grade:12th pass

1 Answers

Arun
3 years ago
Activation energy of forward reaction =0.12−0.02=0.10 kcal
Fraction of molecules which are active or which crossover the energy barrier ( 
A
k
​ 
 )=e 
−E/RT
 
log 
e
​ 
 ( 
A
k
​ 
 )=− 
RT
E
​ 
 
A
k
​ 
 =antilog[ 
2.303RT
−E
​ 
 ]
                 =antilog[ 
2.303×2×325
−0.10×1000
​ 
 ]
                 =antilog[−0.06680]=0.8574
∴ Percentage of reactant molecules crossing over the barrier =0.8574×100=85.74